You recently bought a house with a 10year mortgage loan of TK 125,00,000. The lender charges 12% nominal interest. Mortgage payments are made at the end of each quarter. How much total interest you have to pay (in TK) on this mortgage loan?


1
Expert's answer
2021-12-07T10:00:53-0500

A)Calculation of total interest payable on mortgage:


First we shall calculate quarterly installment amount(P) as follow:


Quarterly interest rate(r)="(12\%\times0.25)=0.03"


No.of payments(n)"=10\times4=40"


Loan amount(A)=TK 12500,000


Quarterly installment amount is:


"=A\frac{r(1+r)^n}{(1+r)^n-1}" ]


"=125,00,000\frac{0.03\times(1+0.03)^{40}}{(1+0.03)^{40}-1}" ]


=TK 540,779.72375


Thus total interest payable is:

Total amount payable on loan-Loan amount

"=(540,779.72375\times40)-125,00,000"


=TK 9131,188.95